OpenStudy (anonymous):

Find the exact value of the expression: cos(sin^-1 1/3 - tan^-1 1/2)

OpenStudy (anonymous):

\[\cos(\sin ^{-1}\frac{ 1 }{ 3 }-\tan ^{-1}\frac{ 1 }{ 2 })\]

OpenStudy (anonymous):

put it all in terms of cos

OpenStudy (anonymous):

How? :/

OpenStudy (anonymous):

let A=\[\sin^{-1} 1/3\] and \[B=\tan^{-1} 1/2\] now sinA=1/3 and tanB=1/2 so cosA=[(2)^3/2]/3..and cosB=[(3)^1/2]/2... so\[A= \cos^{-1} (2^{3/2}) and B=\cos^{-1}(3^{1/2})..\]... now..put AandB in the previous expression...expand it and u'll get the answer..

OpenStudy (anonymous):

cos(x) = sin (x+1.5)

OpenStudy (anonymous):

tan = sin/cos

OpenStudy (anonymous):

also sin x= cos(x-1.5)

OpenStudy (anonymous):

however tan (x) =cos(x-1.5)/cos (x)

OpenStudy (anonymous):

nvrmnd i suppose

OpenStudy (anonymous):